How do you solve by using elimination with 2x-4y=28 4x=17-5y
Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

How do you solve by using elimination with 2x-4y=28 4x=17-5y
=================================================================
we have two equations in two unknowns
==============================================================
2x = 4y+28
divide both sides of = by 2
x = 2y +14
===================================================================
Substitute for x in second equation
===================================================================
4*(2y+14) = 17 -5y
8y +56 = 17 -5y
13y = -39
y = -3
======================
2x -4*(-3) = 28
2x +12 = 28
2x = 16
x = 8
=========================================
x = 8, y = -3

Answer by MathTherapy(10552)


How do you solve by using elimination with 2x-4y=28 4x=17-5y

2x - 4y = 28 ----- eq (i)

4x + 5y = 17 ----- eq (ii)

It's easiest to multiply eq (i) by - 2, or 2 and then add, or subtract both equations to ELIMINATE the x-variable, and determine the y-value. Then substitute the y-value into any of the 2 equations and determine the value of x.

You could also multiply eq (i) by 5 and eq (ii) by 4 to ELIMINATE the y-variable, but this would entail a little bit more work.